OpenStudy (anonymous):

PLEASE HELPPP!!!! A building lot in a city is shaped as a 30-60-90 triangle. The side opposite the 30 angle measures 41 feet. a. Find the length of the side of the lot opposite the 60 angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30 angle in the lot. Write your answers as decimals rounded to four decimal places. Zack is building a model sailboat. The mast will have two inches of height below the base of the main sail. He wants the base of the sail to have a length (I'll draw it out)

OpenStudy (wolfboy):

Connexus acadamy?

OpenStudy (wolfboy):

Test

OpenStudy (wolfboy):

yup i took this two days ago

OpenStudy (anonymous):

Ok so can you help me please??

OpenStudy (wolfboy):

OK i will help

OpenStudy (anonymous):

There is also another one:)

OpenStudy (anonymous):

Thank you:) You are saving my life:)

OpenStudy (anonymous):

Here's the other question: a. What will be the length of the new street? Explain how you found this length. b. The construction cost has been estimated at $140 per linear foot. Estimate the cost for constructing the street. Explain your method for finding the estimate.

OpenStudy (wolfboy):

a. ok for this one we do first what is the \[\sqrt{3}\]

OpenStudy (anonymous):

Is that a square root

OpenStudy (wolfboy):

Yes

OpenStudy (wolfboy):

Step2times the square root by 41

OpenStudy (anonymous):

What square root is that lol o.o

OpenStudy (wolfboy):

Then we have the answer of 71.0141

OpenStudy (wolfboy):

for a

OpenStudy (anonymous):

ok

OpenStudy (wolfboy):

b. 82ft

OpenStudy (wolfboy):

Do you have a scientific couculater?

OpenStudy (anonymous):

ok:) and yes i do

OpenStudy (wolfboy):

ok so enter 41 then click on the button that has sin on it

OpenStudy (anonymous):

-0.158622... right??

OpenStudy (wolfboy):

no its .5

OpenStudy (anonymous):

oh:(

OpenStudy (anonymous):

ok whats next?

OpenStudy (wolfboy):

the cos. and the tan. This you would do the same thing on the couculater so do you just want the answers?

OpenStudy (anonymous):

Yes because i have a deadline til i have to turn in this test

OpenStudy (wolfboy):

ok then cin .66 tan .5773

OpenStudy (anonymous):

which one is this?

OpenStudy (wolfboy):

The rest of question c

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you next one

OpenStudy (wolfboy):

pushypushy lol jk

OpenStudy (anonymous):

im sorry:(

OpenStudy (wolfboy):

its ok but i cannot really help you with the sailboat one though

OpenStudy (anonymous):

oh ok no problem

OpenStudy (wolfboy):

But here was my answer for the street one

OpenStudy (anonymous):

ok

OpenStudy (wolfboy):

a). 4^2 + 9^2 = c 16 + 81 = 97 sqrt97=9.7 b) 9.7* 140=1,358$

OpenStudy (anonymous):

THANK YOU SOOOO MUCH!!!! :D You saved me:D

OpenStudy (wolfboy):

its ok fan me if you like i dont care

OpenStudy (anonymous):

I just did whats your name at connexus??

OpenStudy (wolfboy):

Why?

OpenStudy (anonymous):

just wonering lol :/

OpenStudy (anonymous):

*wondering